OpenStudy (anonymous):

what the common ratio of 80,20,5,5/4... I don't understand how 5 gets to 5/4?!?!?!

OpenStudy (anonymous):

how do you get from 80 to 20?

OpenStudy (anonymous):

-60

OpenStudy (anonymous):

does that same thing work in getting from 20 to 5?

OpenStudy (anonymous):

yes its - 15

OpenStudy (anonymous):

but im not sure if your suppose to simplify 5/4 ? or what? lol

OpenStudy (anonymous):

no, you'd have to subtract 60. think of a different way so that the same operation with the same number will get you from 20 to 5.

OpenStudy (anonymous):

to find the common ratio: \[\frac{ 20 }{ 80 }=\frac{ 5 }{ 20 }=\frac{ \frac{ 5 }{ 4 } }{ 5 }\]

OpenStudy (anonymous):

well I tried doing 20 -60 and it gets me negative 40

OpenStudy (anonymous):

you realise that for all the ratios above, you get 5....so that is the common ratio

OpenStudy (anonymous):

5 is not the common ratio

OpenStudy (anonymous):

@trisha9876 think of another way to get from 80 to 20. not addition or subtraction, but some other way

OpenStudy (anonymous):

@kausarsalley it's not... i'll message you privately

OpenStudy (anonymous):

yea? but in another ratio its 2,4,8,16 and to get from 2 to 4 you have to times it by 2 to get 4 and times 2 x 4 to get 8 and so one an so forth!

OpenStudy (anonymous):

does that make sense? lol

OpenStudy (anonymous):

just think of the different ways to get from 80 to 20, using something other than addition or subtraction. give it a shot

OpenStudy (anonymous):

ok @pgpilot326

OpenStudy (anonymous):

hint... multiplication or division...

OpenStudy (anonymous):

okkaaayyyy I got it!!! thanks guys! and its multiplication! thanks

OpenStudy (anonymous):

can you share the ratio with me?